A rigid container of volume 10 m^3 stores 11.221 k
Answer #1

Consider a rigid container of volume 10 m. Mass water in the container is 11.221 kg. The initial temperature of the water is10 m 11.221 kg = 0.8911 m kg Substitute the data in the above equation as follows: Vi =V+9(V, -VL) 0.8912 = 0.001032 m +9(2.8At 105 C, note the specific volume of saturated vapor and liquid as follows: Specific volume of saturated vapor, V, = 1.4184At 85°C, Specific enthalpy of saturated vapor = 2651.33 kJ Specific enthalpy of saturated liquid = 355.946 Hence, H (85°C) =(At 105°C, Specific enthalpy of saturated vapor = 2683.39 Specific enthalpy of saturated liquid = 440.213 H (105°C)=(1-x)Hz +U = m(U (105°C)-U (85C)) =11.221 kg[1741.23–1027.65]k! = 8007.17 kJ Hence, change internal energy is 8007.17 kJ. Calculate enFrom interpolation, pressure at V, = 0.8911 – m? kg is 1.986 bar. Temperature at saturated vapor is 119.978°C. At 119.978°C,PV-diagram for the above steps is given as follows: P(bar) Critical point 1.986 1.209 0.5784 119.978C -1050 85C Vapor-Liquid